# LayerPano3D: Layered 3D Panorama for Hyper-Immersive Scene Generation

### v1

@@ 3.2 Multi-Layer Panorama Construction
We introduce the Layered 3D Panorama representation based on the following assumption: “an enclosed 3D scene contains a background and various assets positioned in front of it”. In this regard, using Layered 3D Panorama for 3D scene generation is a general approach to handle occlusion for various types of scene. To create a complete scene from the reference panorama, we propose to decompose it into $N+1$ layers along the depth dimension. As shown in Figure 2, these layers, arranged from farthest to nearest, represent both the scene background (layer 0) and the layouts situated behind the observation point.
@@
Layer Decomposition. As shown in Figure 2, the reference panorama is decomposed by first identifying the scene assets and then cluster these assets in different layers according to depth. First, we employ an off-the-shelf panoptic segmentation model [45] pretrained on ADE20K [46] to automatically find all scene assets visible in the reference panorama. A good layer decomposition requires that the layer assets share a similar depth level within layers and are distant from assets in other layers. In this sense, we assign each asset a depth value and apply K-Means to cluster these masks into different groups. Given the reference panorama depth map, the depth value for each asset mask is determined by calculating the 75th percentile of the depth values within the masked region. According to the depth values, the assets are clustered into $N$ groups from layer 00 to $N-1$ and are merged into layer masks to guide the subsequent layer completion.
@@
Layer Completion. With the layer mask, we focus on completing the unseen content caused by asset occlusion. In order to synthesize background pixels instead of creating new elements, we finetune PanFusion [11] with the ERP-perspective aware cross-attention as the panoramic background inpainter. Specifically, at each layer, our model takes the layer mask $M_{l}$, the reference panorama, and the “empty scene, nothing” [47] prompt as input, and output coherent content at the masked area. The inpainted panorama at layer $l$ is denoted $P_{l}$ and is used as supervision to the subsequent panoramic 3D Gaussian scene optimization. Note that, we additionally apply SAM [48] to extend the layer mask, based on the inpainted panorama from the previous layer, to eliminate unwanted new generations from inpainting.
@@
Moreover, to enable free rendering in 3D, where observers can examine scenes from varying distances, the unprocessed textures of distant assets may appear blurred as the observer approaches. Therefore, distant layers require higher resolution to preserve texture details at different viewpoints. To address this, Super Resolution (SR) module [49] is employed to enhance the resolution of the layered panorama from layer 00 (background layer) to layer $N$ (reference panorama), achieving a 2×2$\times$2 × upscale in resolution. SR processing significantly improves the texture quality of distant objects, maintaining their visual clarity and texture details even when observed from a closer perspective.
Layer Alignment. Given the Layered 3D RGB Panorama $[P_{l}]_{l=0}^{N}$, we perform the depth prediction and alignment to ensure consistency in a shared space. To begin with, we apply the 360MonoDepth [50], to first estimate the layer $N$ (reference panorama) as the reference depth $P_{depth}^{N}$. The predicted depth $P_{depth}$ inherently contains extremely small values from normalization, causing peak-like artifacts under the reference camera position. Consequently, we post-process the directly output disparity map $\hat{P}_{disp}$ with the following transformation to mitigate the geometry distortion:
$D_{\text{bias}}$ $=\left(\max(\hat{P}_{\text{disp}})-\min(\hat{P}_{\text{disp}})\right)\ /\ 4,$ (4)
$\hat{P}_{\text{depth}}$ $=\max(\hat{P}_{\text{disp}})-\hat{P}_{\text{disp}}+D_{\text{bias}},$
$P_{\text{depth}}$ $=\hat{P}_{\text{depth}}\ /\ \max(\hat{P}_{\text{depth}}).$
@@
To align the layer depth in 3D space, we find it infeasible to simply compute a global shift and scale as in  [4, 6] due to the nonlinear nature of ERP. Therefore, we leverage depth inpainting model $\mathcal{F}_{depth}$ from [51] to directly restore depth values based on the inpainted RGB pixels. $\mathcal{F}_{depth}$ harnesses strong generalizability from large-scale diffusion prior and synthesizes inpainted depth values at an aligned scale with the base depth. We start from reference panoramic depth $P_{depth}^{N}$ to implement step-by-step restoration from layer $N-1$ to layer 00:
$P_{depth}^{l}=\mathcal{F}_{depth}(P_{l},M_{l}\odot P_{depth}^{l+1}),$ (5)
@@
3.3 Panoramic 3D Gaussian Scene Optimization
Figure 4: Illustration of the Gaussian Selector. Given the new asset point cloud, the Gaussian Selector identifies the active Gaussians for next layer’s optimization. (a) Optimized Gaussians at layer $l$ partially blocks new asset point cloud in layer $l+1$. (b) By computing the viewing directions and the distance to camera, Gaussian Selector identifies the blocking Gaussians. (c) The marked Gaussians are labeled as active, and merged with the new points for optimization. (d) Showed a special case where the optimized Gaussian center is behind the new points, but also identifies as an active Gaussian because its volume scale overlaps with the points (case 1).
3D Scene Initialization. To enable free 3D exploration, we lift the Layered 3D Panorama to 3D Gaussians [19], where the Gaussians are initialized from the layered 3D panoramic point clouds. Considering the intrinsic spherical structure of panorama, we can easily transform an equirectangular image $P\in\mathbb{R}^{H\times W\times 3}$ into 3D point cloud $S(\theta,\phi,P_{depth})$. Each pixel $(u,v)$ is represented as a 3D point and the angles $\theta,\phi$ are computed as $\theta=(2u/W-1)\pi$, $\phi=(2v/H-1)\pi/2$.
@@ 3.3 Panoramic 3D Gaussian Scene Optimization
$X$ $=P_{depth}(\theta_{u},\phi_{v})\cos{\phi_{v}}\cos{\theta_{u}},$ (6)
@@
3D Scene Refinement. During scene refinement, we devise two types of Gaussian training schemes for varying scene content: the base Gaussian for reconstructing the scene background and the layer Gaussian for optimizing scene layouts. Additionally, a Gaussian selector module is introduced between layer Gaussians to facilitate scene composition.
@@
We observe that the quality of the optimized scene is easily hampered by unaligned layers, and sometimes $\mathcal{F}_{depth}$ does fail to produce perfectly aligned layer depths. Gaussians at layer $l$ could span into unwanted depth levels and block assets in the subsequent layer, as illustrated in Figure 4(a). To handle this issue, we introduce the Gaussian selector module to detect these conflicted Gaussians, re-activate them from frozen, and optimize them away from the blockage. First, the selector computes the distance vector from the camera center $\mathbf{o}=(0,0,0)$ to each new point $\mathbf{p}$, as in Figure 4(b). The absolute distance from asset points $\mathbf{p}$ and scene Gaussians $\mathbf{g}$ to the camera is denoted as $d_{\mathbf{p}}$ and $d_{\mathbf{g}}$ respectively:
$d_{\mathbf{p}}=||\mathbf{p}-\mathbf{o}||_{2},\quad d_{\mathbf{g}}=||\mathbf{g}-\mathbf{o}||_{2}.$ (7)
By examining all Gaussians that on the same ray with asset points but at a closer distance: $\mathbf{p}\ /\ d_{\mathbf{p}}=\mathbf{g}\ /\ d_{\mathbf{g}},\quad d_{\mathbf{g}}<d_{\mathbf{p}},$ we mark them as active (Figure 4(c)). We also observe that despite the Gaussian center being behind the asset points, if the scaling of the Gaussian covers the point, then the Gaussian is still affecting the overall rendering quality. Hence, we update the gaussian distance as
$d^{\prime}_{\mathbf{g}}=d^{\prime}_{\mathbf{g}}-\max\{\mathbf{s}(\mathbf{g})\},$ (8)
where $\mathbf{s}$ is the scaling of the Gaussian $\mathbf{g}$. If $d_{\mathbf{g_{1}}}>d_{\mathbf{p_{1}}}$ and $d^{\prime}_{\mathbf{g_{1}}}<d_{\mathbf{p_{1}}}$ as in Figure 4(d), the referred Gaussian is also marked active. For efficient memory storage and fast look-up, we hash the distance vectors into a 3D grid. The mapping function from vector coordinates to grid indices writes: $f(\mathbf{p})=\text{ceil}(\beta_{3}\log(\textbf{p}+1))$.
Figure 5: Qualitative comparisons in Panorama Generation. We compare 2D panorama generation methods across three dimensions: Creativity (red), Resolution (green), and Spherical Structure (blue). LayerPano3D produces high-resolution, creative, and consistent panoramas while accommodating multi-style generation requirements.
4 Experiments
4.1 Implementation Details
In the anchor view synthesis stage, we train the model starting from SDXL-base-1.0 [18] with a batch size of 16 and learning rate of $10^{-5}$ for only 10K iterations on the 11k data pairs. The training is done using 4 NVIDIA A100 GPUs for 7 hours. For inference, we initialize four $512\times 512$ grids with random colored noise, subsequently superimposing a white cross-line with a pixel width of 10 to ensure that the final inference result can be accurately segmented into four grids of size $507\times 507$. In the reference panorama synthesis stage, we leverage the Stable Diffusion Inpainting pipeline [41] from huggingface [52] as the inpainter and repurposed Diffusion360 [43] as the outpainter to synthesize $360^{\circ}\times 180^{\circ}$ FOV panorama.
In the layered panorama construction stage, we employ OneFormer [45] to obtain the panoptic segmentation map for the reference panorama. Background categories are manually determined (i.e. sky, floor, ceiling, etc.) to filter out background components in asset masks. Generally, we cluster all asset masks into $N=3$ layers via KNN and merge all masks within each layer to form a unified layer mask. With the obtained layer mask, we combine our fine-tuned panorama inpainting model with LaMa [42] to achieve multi-layer completion and apply 360MonoDepth [50] to predict the reference panorama depth.
@@ 4.1 Implementation Details
In the 3D panoramic scene optimization stage, we lift the panorama RGBD into 3D point clouds. For scene initialization, we set $\beta_{1}$ to $0.0001$ and $\beta_{2}$ to $4$ based on empirical practice. These point clouds are used to initialize the base Gaussian model and the layer Gaussian model. During the scene refinement stage, we optimize the base Gaussian model for 3,000 iterations, then the layer Gaussians each for 2,000 iterations. The training objective for base and layer Gaussian is the L1 loss and D-SSIM term between the ground-truth views and the rendered views. We use a single 80G A100 GPU for reconstruction and the reconstruction time for each layer is 1.5 minutes on average for $1024\times 1024$ resolution inputs.
4.2 Comparison Methods.
To evaluate the performance of our approach in the context of text-driven 3D panoramic scene generation, We compare with previous methods in two phases: 2D Panorama Generation and 3D Panoramic Scene Reconstruction. For 2D Panorama Generation, we choose four approaches to check the quality and creativity of 2D panorama: Text2light [15] adopts the VQGAN [53] structure to synthesize an HDR panorama image from text in a two-stage auto-regressive manner. Panfusion [11] designs a novel dual-branch diffusion model to generate a 360-degree image from a text prompt. Diffusion360 [43] directly uses the DreamBooth [54] training on the panorama dataset [55]. LDM3D-pano [12] extends LDM3D [56] to generate a realistic RGB panorama and its corresponding depth. For 3D Panoramic Scene Reconstruction, we select three methods to compare the 3D performance of scenes: LucidDreamer [4] starts from a single image and a text prompt to create a 3D-GS scene. Text2Room [6] generates a textured 3D mesh scene with preset camera trajectory. Dreamscene360 [16] creates panoramic 3DGS scene with full $360^{\circ}$ coverage from text prompts. Note that, since DreamScene360 does not open-source its official code, we reproduce its model for comparison.
4.3 Qualitative Comparison
@@
2D Panorama Generation. We show some qualitative comparisons with several state-of-the-art panorama generation works in Figure 5. Text2light ignores style prompts due to being trained on a realistic HDRI dataset based on the VQGAN structure, and the components in the generated panorama are relatively simple. The results by PanFusion are ambiguous and low in quality, while it can adapt to style changes due to its operation in panorama and perspective domains. LDM3D-pano is insensitive to style prompts and the results are low-resolution. The instances generated by Diffusion360 appear to be of higher quality compared to the above, but do not have the inherent spherical structural property of panoramas, and the style seems to be the same. This is because the model is merely trained on the panorama dataset based on Dreambooth [54] without injecting any prior knowledge. Our method achieves the highest quality both in consistency and diversity, presenting creative and reasonable generations.
@@
3D Panoramic Scene Reconstruction. To more precisely evaluate the quality of 3D panoramic scene reconstruction, we present qualitative comparisons with Text2Room [6], LucidDreamer [4], and DreamScene360 [16] across two dimensions. For full $360^{\circ}\times 180^{\circ}$ view consistency, we input an image with a text prompt to guide scene creation and render multiple views from the center of the scene. This approach synthesizes the texture of the panoramic environment, allowing for a more intuitive assessment after generation. As shown in Figure 6, LucidDreamer [4] and Text2room [6] fail to cover the full $360^{\circ}\times 180^{\circ}$ view, resulting in semantic incoherence and artifacts due to their successive inpainting-based strategy. DreamScene360 [16] supports a $360^{\circ}\times 180^{\circ}$ view at a single fixed viewpoint, but it deviates from the input conditions, and the quality of the generated results is relatively low. In contrast, our model excels in maintaining full $360^{\circ}\times 180^{\circ}$ view consistency while demonstrating superior content creativity. To evaluate the capability for free trajectory rendering, we design a zigzag trajectory to guide the camera’s movement through the scene, with novel view renderings sampled along the trajectory for comparison. As illustrated in Figure 7, we showed 3 random samples from this fixed flythrough trajectory. Compared with the other three methods, our model achieves a more complete 3D scene with consistent textures and a reasonable geometric structure.
@@
Figure 6: Qualitative comparisons in full 360°×180° Scene. We compare the panorama and multiple views of the scene generated by four methods.  LayerPano3D  exhibits more consistent content with full $360^{\circ}\times 180^{\circ}$ coverage.
@@
Figure 7: Qualitative comparisons in Scene Free Exploration. We show the novel view renderings along a zigzag trajectory to compare the capability of free trajectory rendering. Our method is able to maintain high-quality content rendering and does not show distortion or gaps in unseen space, which shows the ability of  LayerPano3D  to create hyper-immersive panoramic scenes.
Figure 8: Analysis on the Layer Completion Inpainting. We present the panorama inpainting results for three methods guided by the same text prompt: “empty scene, nothing” [47]. Our model effectively handles complex scenarios, delivering clear results with consistent and coherent structures.
4.4 Quantitative Comparison
2D Panorama Generation. We adopt three metrics for quantitative comparisons: 1) Intra-Style [57, 34] evaluates the coherence of the generated panoramas; 2) FID [58] evaluates both fidelity and diversity; 3) CLIP [59] measures the compatibility of results and input prompts. We selected 73 prompts to generate panoramas for each method and averaged the results for all metrics. Moreover, a user study is also conducted to further evaluate the quality of panoramas, where we project 4 views at a fixed FOV ($90^{\circ}$) to the user for sorting. As shown in Table I, our method achieves the best scores in both FID and CLIP metrics, indicating its high fidelity and strong compatibility with the input prompts. The Intra-Style results demonstrate that our model achieves global coherence across the image, consistently maintaining the overall style. While Text2light has a smaller Intra-Style score, this is primarily due to its tendency to generate monotonous results with extensive uniform color block backgrounds. We also attained outstanding results in aesthetic metrics, marginally trailing Diffusion360. However, as discussed in Section 4.3, panoramas generated by Diffusion360 lack the inherent spherical structural property, resulting in an unreasonable final perspective projection.
Moreover, we conducted a user study comparing generated panoramas, evaluating them across three dimensions: Coherence (internal consistency), Plausibility (plausibility of perspective views), and Compatibility (align with prompts). Following [60], we use the Average User Ranking (AUR) as a preference metric where users rank each result on a scale of 1 to 5 (lower is worse). We invite 29 users including graduate students that expertise in 3D and average users to rank the 60 results from 5 methods individually. The results in Table I demonstrate that our method outperforms others by a significant margin across all dimensions.
@@ 4.4 Quantitative Comparison
TABLE I: Quantitative comparison with SoTA methods on 2D Panorama Generation. Bold indicates the best result, and underline indicates the second-best result.
      Method       FID $\downarrow$       Intra-Style $\downarrow$       CLIP $\uparrow$       User Study (AUR)
      Coherence $\uparrow$       Plausibility $\uparrow$       Compatibility $\uparrow$
      Text2light [15]       286.90       0.31       18.69       2.95       2.47       1.89
      Panfusion [11]       283.80       18.66       21.22       3.06       3.03       3.26
      Diffusion360 [43]       274.03       3.70       21.65       3.08       3.22       3.24
      LDM3D-pano [12]       298.64       8.05       20.58       2.56       2.84       2.87
      Ours       255.60       2.24       22.08       3.35       3.44       3.73
TABLE II: Qualitative comparison with SoTA methods on 3D Panoramic Scene. Bold indicates the best result.
@@
NIQE $\downarrow$ BRISQUE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Roll-Mean $\downarrow$ Roll-Var $\downarrow$
@@
Text2room [6] 5.117 44.598 30.437 0.890 0.033 1.877 1.566 1.76 2.28
LucidDreamer [4] 5.877 54.230 31.020 0.973 0.029 2.775 2.167 1.85 1.29
DreamScene360 [16] 5.083 39.403 28.925 0.941 0.050 1.696 2.524 2.94 2.90
Ours 4.374 39.006 40.797 0.980 0.012 0.540 0.028 3.45 3.53

### v2

@@ 3.1. Multi-Layer Panorama Generation
We introduce the Layered 3D Panorama representation based on the following assumption: “an enclosed 3D scene contains a background and various assets positioned in front of it”. In this regard, using Layered 3D Panorama for 3D scene generation is a general approach to handle occlusion for various types of scene. To create a complete scene, we first generate a high-quality reference panorama from a single text prompt and decompose it into $N+1$ layers along the depth dimension. As shown in Fig. 2, these layers, arranged from the farthest to the nearest, represent both the scene background (layer 0) and the layouts behind the observation point.
3.1.1. Reference Panorama Generation and Upright360 .
A good panorama captures rich scene details, creating an immersive and comprehensive view. This depth of detail fosters a stronger connection and deeper comprehension of the scene. Unlike the advancement in standard image generation, panorama generation still faces quality gaps. To ensure upright and geometrically consistent scenes, we curate a new dataset, namely Upright360 , comprising high quality, upright panorama images, and finetune the Flux (Labs, 2023) with LoRA on the Upright360 dataset. This lightweight training approach achieves optimal performance even with limited high-quality panorama data and can be directly extended to subsequent tasks, for example panorama inpainting for layer completion. For data curation, we first collect around 15k raw panorama images: 9684 panorama image samples are collected from Matterport3D (Chang et al., 2017), 1824 images from the web, and 3592 synthetic panoramas generated from Blockadelabs. Based on this, we leverage GeoCalib (Veicht et al., 2024), a state-of-the-art single-perspective-image calibration method, to filter upright panoramas. Specifically, we employ the Equirectangular Projection (ERP), a mapping technique that projects a 3D sphere onto a 2D plane, to generate four perspective views from each panorama. These views are extracted at a fixed field of view (FOV) of 90°, an elevation of 0°, and four distinct azimuths (0°, 90°, 180°, 270°). We then use GeoCalib to calibrate the four views, computing their pitch and roll variances for filtering. Panoramas with variances exceeding 1.0 are classified as non-upright and excluded: $Var({pitch_{1:4}})>1.0,Var({roll_{1:4}})>1.0$, resulting in the creation of the final Upright360 dataset. This dataset comprises 9423 high-quality panoramas, rigorously filtered from the original collection. Building upon the Upright360 dataset, we fine-tune Flux to develop a panorama LoRA (Hu et al., 2021b) for reference panorama generation.
3.1.2. Layer Decomposition.
@@
As shown in Fig. 2, the reference panorama is decomposed by first identifying the scene assets and then cluster these assets in different layers according to depth. First, we employ an off-the-shelf panoptic segmentation model (Jain et al., 2023) pretrained on ADE20K (Zhou et al., 2017) to automatically find all scene assets visible in the reference panorama. A good layer decomposition requires that the layer assets share a similar depth level within layers and are distant from assets in other layers. In this sense, we assign each asset a depth value and apply K-Means to cluster these masks into different groups. Given the reference panorama depth map, the depth value for each asset mask is determined by calculating the 75th percentile of the depth values within the masked region. According to the depth values, the assets are clustered into $N$ groups from layer 00 to $N-1$ and are merged into layer masks to guide the subsequent layer completion.
3.1.3. Layer Completion.
@@
With the layer mask, we focus on completing the unseen content caused by asset occlusion. In order to synthesize background pixels instead of creating new elements, we directly utilize the above trained panorama lora and integrate it into the Flux-Fill model to accomplish domain transfer, thereby employing it as a panoramic canvas inpainter. Specifically, at each layer, our model takes the layer mask $M_{l}$, the reference panorama, and the “empty scene, nothing” (Zhang and Agrawala, 2024) prompt as input, and output coherent content at the masked area. The inpainted panorama at layer $l$ is denoted $P_{l}$ and is used as supervision to the subsequent panoramic 3D Gaussian scene optimization. Note that, we additionally apply SAM (Kirillov et al., 2023) to extend the layer mask, based on the inpainted panorama from the previous layer, to eliminate unwanted new generations from inpainting.
@@
Moreover, to enable large-range rendering in 3D, where observers can examine scenes from varying distances, the unprocessed textures of distant assets may appear blurred as the observer approaches. Therefore, distant layers require higher resolution to preserve texture details at different viewpoints. To address this, Super Resolution (SR) module (Yang et al., 2023) is employed to enhance the resolution of the layered panorama from layer 00 (background layer) to layer $N$ (reference panorama), achieving a 2×2$\times$2 × upscale in resolution. SR processing significantly improves the texture quality of distant objects, maintaining their visual clarity and texture details even when observed from a closer perspective.
3.1.4. Layer Alignment.
@@
Given the Layered 3D RGB Panorama $[P_{l}]_{l=0}^{N}$, we perform the depth prediction and alignment to ensure consistency in a shared space. To begin with, we apply the 360MonoDepth (Rey-Area et al., 2022), to first estimate the layer $N$ (reference panorama) as the reference depth $P_{depth}^{N}$. Then, to align the layer depth in 3D space, we find it infeasible to simply compute a global shift and scale as in  (Chung et al., 2023; Höllein et al., 2023) due to the nonlinear nature of ERP. Therefore, we leverage depth inpainting model $\mathcal{F}_{depth}$ from (Liu et al., 2024) to directly restore depth values based on the inpainted RGB pixels. $\mathcal{F}_{depth}$ harnesses strong generalizability from large-scale diffusion prior and synthesizes inpainted depth values at an aligned scale with the base depth. We start from reference panoramic depth $P_{depth}^{N}$ to implement step-by-step restoration from layer $N-1$ to layer 00:
(1) $P_{depth}^{l}=\mathcal{F}_{depth}(P_{l},M_{l}\odot P_{depth}^{l+1}),$
@@
3.2. Panoramic 3D Gaussian Scene Optimization
Figure 3. Illustration of the Gaussian Selector. Given the new asset point cloud, the Gaussian Selector identifies the active Gaussians for next layer’s optimization.
3.2.1. 3D Scene Initialization.
To enable large-range 3D exploration, we lift the Layered 3D Panorama to 3D Gaussians (Kerbl et al., 2023), where the Gaussians are initialized from the layered 3D panoramic point clouds. Considering the intrinsic spherical structure of panorama, we can easily transform an equirectangular image $P\in\mathbb{R}^{H\times W\times 3}$ into 3D point cloud $S(\theta,\phi,P_{depth})$. Each pixel $(u,v)$ is represented as a 3D point and the angles $\theta,\phi$ are computed as $\theta=(2u/W-1)\pi$, $\phi=(2v/H-1)\pi/2$.
@@ 3.2. Panoramic 3D Gaussian Scene Optimization
(2) $X$ $=P_{depth}(\theta_{u},\phi_{v})\cos{\phi_{v}}\cos{\theta_{u}},$
@@
3.2.2. 3D Scene Refinement.
During scene refinement, we devise two types of Gaussian training schemes for varying scene content: the base Gaussian for reconstructing the scene background and the layer Gaussian for optimizing scene layouts. Additionally, a Gaussian selector module is introduced between layer Gaussians to facilitate scene composition.
@@
We observe that the quality of the optimized scene is easily hampered by unaligned layers, and sometimes $\mathcal{F}_{depth}$ does fail to produce perfectly aligned layer depths. Gaussians at layer $l$ could span into unwanted depth levels and block assets in the subsequent layer, as illustrated in Fig. 3(a). To handle this issue, we introduce the Gaussian selector module to detect these conflicted Gaussians, re-activate them from frozen, and optimize them away from the blockage. First, the selector computes the distance vector from the camera center $\mathbf{o}=(0,0,0)$ to each new point $\mathbf{p}$, as in Fig. 3(b). The absolute distance from asset points $\mathbf{p}$ and scene Gaussians $\mathbf{g}$ to the camera is denoted as $d_{\mathbf{p}}$ and $d_{\mathbf{g}}$ respectively:$d_{\mathbf{p}}=||\mathbf{p}-\mathbf{o}||_{2},\quad d_{\mathbf{g}}=||\mathbf{g}-\mathbf{o}||_{2}$. By examining all Gaussians that on the same ray with asset points but at a closer distance: $\mathbf{p}\ /\ d_{\mathbf{p}}=\mathbf{g}\ /\ d_{\mathbf{g}},\quad d_{\mathbf{g}}<d_{\mathbf{p}},$ we mark them as active ( Fig. 3(c)). For efficient memory storage and fast look-up, we hash the distance vectors into a 3D grid. The mapping function from vector coordinates to grid indices writes: $f(\mathbf{p})=\text{ceil}(\beta_{3}\log(\textbf{p}+1))$.
4. Experiments
4.1. Implementation Details
@@
In the layered panorama construction stage, we train the panorama LoRA starting from Flux (Labs, 2023) with a batch size of 1 and learning rate of $10^{-5}$ for 100K iterations on our curated Upright360 dataset. The training is done using 6 NVIDIA A100 GPUs for 3 days. For Layer Decomposition, we employ OneFormer (Jain et al., 2023) to obtain the panoptic segmentation map for the reference panorama. Background categories are manually determined (i.e. sky, floor, ceiling, etc.) to filter out background components in asset masks. Generally, we cluster all asset masks into $N=3$ layers via KNN and merge all masks within each layer to form a unified layer mask. With the obtained layer mask, we integrate the above trained panorama LoRA into FLux-Fill model and combine with LaMa (Suvorov et al., 2021) to achieve multi-layer completion and apply 360MonoDepth (Rey-Area et al., 2022) to predict the reference panorama depth. In the 3D panoramic scene optimization stage, we lift the panorama RGBD into 3D point clouds. For scene initialization, we set $\beta_{1}$ to $0.0001$ and $\beta_{2}$ to $4$ based on empirical practice. These point clouds are used to initialize the base Gaussian model and the layer Gaussian model. During the scene refinement stage, we optimize the base Gaussian model for 3,000 iterations, then the layer Gaussians each for 2,000 iterations. The training objective for base and layer Gaussian is the L1 loss and D-SSIM term between the ground-truth views and the rendered views. We use a single 80G A100 GPU for reconstruction and the reconstruction time for each layer is 1.5 minutes on average for $1024\times 1024$ resolution inputs.
4.2. Comparison Methods.
To evaluate the performance of our approach in text-driven 3D panoramic scene generation domain, we compare with existing methods in two phases: 2D Panorama Generation and 3D Panoramic Scene Reconstruction. For 2D Panorama Generation, we compare the quality and creativity of 2D panorama with Text2light (Chen et al., 2022) (GAN-based HDR panorama generation), Diffusion360 (Feng et al., 2023) (diffusion-based text-to-panorama generation) and Panfusion (Zhang et al., 2024) (dual-branch diffusion-based generation). For 3D Panoramic Scene Reconstruction, we compare with Text2Room (Höllein et al., 2023) (iterative indoor scene expansion with textured mesh), LucidDreamer (Chung et al., 2023) (single-view scene generation with 3DGS), and Dreamscene360 (Zhou et al., 2024b) (text-guided panoramic 3DGS scene generation).
4.3. Qualitative Comparison
4.3.1. 2D Panorama Generation.
@@ 4.3. Qualitative Comparison
We show some qualitative comparisons with several state-of-the-art panorama generation works in Fig. 6. Text2Light (Chen et al., 2022) struggles to effectively interpret text prompt due to being trained on a realistic HDRI dataset based on the VQGAN structure, and the components in the generated panorama are relatively simple. The results by PanFusion (Zhang et al., 2024) are ambiguous and low in quality. While the instances generated by Diffusion360 (Feng et al., 2023) exhibit superior quality in comparison to the aforementioned methods, they lack intricate scene details and are prone to the generation of artifacts. In contrast, our method achieves the highest quality, presenting creative and reasonable generations.
4.3.2. 3D Panoramic Scene Reconstruction.
@@
We present qualitative comparisons with Text2Room (Höllein et al., 2023), LucidDreamer (Chung et al., 2023), and DreamScene360 (Zhou et al., 2024b) across two dimensions. First, for full $360^{\circ}\times 180^{\circ}$ view consistency, we render multiple views from scene center point with single input image and text prompts. As shown in Fig. 4, LucidDreamer (Chung et al., 2023) and Text2room (Höllein et al., 2023) fail to cover the full $360^{\circ}\times 180^{\circ}$ view, resulting in semantic incoherence and artifacts due to their successive inpainting-based strategy. DreamScene360 (Zhou et al., 2024b) supports a $360^{\circ}\times 180^{\circ}$ view at a single fixed viewpoint, but the quality of the generated results is relatively low. In contrast, our model excels in maintaining full $360^{\circ}\times 180^{\circ}$ view consistency while demonstrating superior content creativity. Second, to evaluate novel path rendering, we design a zigzag trajectory to guide the camera’s movement through the scene, with novel view renderings sampled along the trajectory for comparison. Fig. 5 shows 6 random samples from this fixed flythrough trajectory. Compared with all three methods, our model achieves a more complete 3D scene with consistent textures and a reasonable geometric structure.
@@
Figure 4. Qualitative comparisons in full 360°×180° Scene. We compare the panorama and multiple views of the scene generated by four methods.  LayerPano3D  exhibits consistent and rich details across full $360^{\circ}\times 180^{\circ}$ coverage, while other methods show obvious inconsistencies and disorganized patterns in regions that deviate from the input view.
@@
Figure 5. Qualitative comparisons in Large-range Scene Exploration. We show the novel view renderings along a zigzag trajectory to compare the capability of large-range scene exploration. Our method is able to maintain high-quality content rendering and does not show distortion or gaps in unseen space, which shows the ability of  LayerPano3D  to create hyper-immersive panoramic scenes.
@@
Table 1. Quantitative comparison with SoTA methods on 2D Panorama Generation. Bold indicates the best result.
Method FID $\downarrow$ Aesthetic $\uparrow$ CLIP $\uparrow$ User Study (AUR) $\uparrow$
Text2light 286.90 4.57 18.69 1.34
Panfusion 283.80 4.78 21.22 2.38
Diffusion360 274.03 5.07 21.65 2.52
Ours 223.51 5.86 22.25 3.76
Table 2. Qualitative comparison with SoTA methods on 3D Panoramic Scene.  LayerPano3D  achieves high-quality reconstruction and novel view synthesis while maintaining upright panoramic scene compared to other methods.
@@
NIQE $\downarrow$ BRISQUE $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Pitch-Mean $\downarrow$ Pitch-Var $\downarrow$
@@
Text2room (Höllein et al., 2023) 5.231 46.127 30.126 0.882 0.038 2.029 1.724 1.69 2.31
LucidDreamer (Chung et al., 2023) 5.822 52.102 36.108 0.954 0.026 2.813 2.189 1.81 1.31
DreamScene360 (Zhou et al., 2024b) 5.051 39.891 30.056 0.958 0.062 1.328 2.018 2.86 2.77
Ours 4.023 38.287 42.057 0.986 0.015 0.732 0.032 3.64 3.61
4.4. Quantitative Comparison
4.4.1. 2D Panorama Generation.
We adopt three metrics for quantitative comparisons: 1) FID (Heusel et al., 2017) evaluates both fidelity and diversity; 2) Aesthetic (Schuhmann et al., 2022) evaluates the aesthetics of panorama; 3) CLIP (Hessel et al., 2021) measures the compatibility of results with input prompts. Moreover, a user study is also conducted to further evaluate the quality of panoramas, where we project 4 views at a fixed FOV ($90^{\circ}$) to the user for sorting. Here, we report the average Average User Ranking (AUR) (Zhang et al., 2023a), which is computed based on an integrated assessment of coherence, plausibility, aesthetics, and compatibility dimensions. As shown in Tab. 1, our method achieves the best scores among all quantitative metrics and human evaluation, demonstrating its fidelity, alignment with text and overall consistency.
4.4.2. 3D Panoramic Scene Reconstruction.